Auto pistons at Wemming​ Chung's plant in Shanghai are produced in a forging​ process, and the diameter is a critical factor that must be controlled. From sample sizes of

1010

pistons produced each​ day, the mean and the range of this diameter have been as​ follows:

                                                               

Day

Mean Bold x overbarx

​(mm)

Range R

​(mm)

1

156.9156.9

4.24.2

2

153.2153.2

4.64.6

3

153.6153.6

4.14.1

4

155.5155.5

5.05.0

5

156.6156.6

4.54.5

​a) What is the value of

x double overbarx​?

x double overbarxequals=nothing

mm ​(round your response to two decimal​ places).

​b) What is the value of

Upper R overbarR​?

Upper R overbarRequals=nothing

mm ​(round your response to two decimal​ places).

​c) What are the

UCL Subscript x overbarUCLx

and

LCL Subscript x overbarLCLx

using

​3-sigma​?

Upper Control Limit

​(UCL Subscript x overbarUCLx​)

​=

nothing

mm ​(round your response to two decimal​ places).

Lower Control Limit

​(LCL Subscript x overbarLCLx​)

​=

nothing

mm ​(round your response to two decimal​ places).

​d) What are the

UCL Subscript Upper RUCLR

and

LCL Subscript Upper RLCLR

using

​3-sigma​?

Upper Control Limit

​(UCL Subscript Upper RUCLR​)

​=

nothing

mm ​(round your response to two decimal​ places).

Lower Control Limit

​(LCL Subscript Upper RLCLR​)

​=

nothing

mm ​(round your response to two decimal​ places).

​e) If the true diameter mean should be 155 mm and you want this as your center​ (nominal) line, what are the new

UCL Subscript x overbarUCLx

and

LCL Subscript x overbarLCLx​?

Upper Control Limit

​(UCL Subscript x overbarUCLx​)

​=

nothing

mm ​(round your response to two decimal​ places).

Lower Control Limit

​(LCL Subscript x overbarLCLx​)

​=

nothing

mm ​(round your response to two decimal​ places).
Answer #1

Answer a and b

R= Max- Min X-bar= average of sample
4.20 156.90
4.60 153.20
4.10 153.60
5.00 155.50
4.50 156.10
4.48 155.06
R-bar (Average of above values) X-bar-bar (average of above values)

Answer c: D4, D3 and A2 are taken from table of factors computing 3 sigma control limits, sample size=10

A2 (n=10) 0.308
control limits for X bar,
CL or Xbarbar 155.06
UCL=Xbarbar+ (A2)*Rbar 156.44
LCL=Xbarbar- (A2)*Rbar 153.68

Answer d:

D4 (n=10)= 1.777
D3 (n=10)= 0.223
control limits for Range,
CL or R-bar= 4.48
UCL=R-bar*D4 7.96
LCL=R-bar*D3 1.00

Note: Chegg's guidelines mandate to solve one question or 1st four subparts of a question at a time. Please post again for remaining questions.